Is (1,-3) a solution of y=-2x-1?
Found 2 solutions by rapaljer, lenny460:
Answer by rapaljer(4671)

The point (1,-3) means to substitute x=1, and see if y=-3.

y=-2x-1
y=-2(1)-1
y=-2-1=-3. YES!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida

Answer by lenny460(1073)

Draw the straight line graph of:
y -2x - 1
From the above we get the following coordinates:
(2,-5)
(1,-3)
(0,-1)
(-1,1)
(-2,3)
Yes coordinate (1,-3) is a solution in the
straight line graph of y = -2x - 1
The Slope (m) rise/run = -4/2 = -2
The Slope (m) = -2
the y-intercept is: -1


obuong3@aol.com